OpenStudy (anonymous):

what is (1+i)/i(2+3i) ? how do I do it?

OpenStudy (solomonzelman):

1) multiply top and bottom times i. 2) use conjugate (multiply top and bottom times `2-3i`

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ \frac{(1+i)}{i(2+3i)} }\) \(\LARGE\color{black}{ \frac{(1+i)\color{darkgoldenrod}{\times i} }{i(2+3i)\color{darkgoldenrod}{\times i}} }\) \(\LARGE\color{black}{ \frac{1i-1}{-1(2+3i)} }\) \(\LARGE\color{black}{ \frac{i-1}{-1(2+3i)} }\) \(\LARGE\color{black}{ \frac{-(-i+1)}{-1(2+3i)} }\) \(\LARGE\color{black}{ \frac{-(1+i)}{-1(2+3i)} }\) \(\LARGE\color{black}{ \frac{1+i}{2+3i} }\)

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ \frac{(1+i)\color{red}{\times(2-3i)}}{(2+3i)\color{red}{\times(2-3i)}} }\) applying `(a+bi)(a-bi)=a²+b²` to the bottom \(\LARGE\color{black}{ \frac{(1+i)\color{red}{\times(2-3i)}}{4+3} }\) \(\LARGE\color{black}{ \frac{(1+i)\color{red}{\times(2-3i)}}{7} }\)

OpenStudy (solomonzelman):

Then, \(\LARGE\color{black}{ \frac{2-3i+2i+3}{7} }\) \(\LARGE\color{black}{ \frac{5-i}{7} }\)

OpenStudy (solomonzelman):

Any questions ?

OpenStudy (anonymous):

thankyou so much! made pretty good sense to me :))